OpenStudy (anonymous):

Simplify: \[\frac{ \cot ^{2}x }{ 1+cscx }+\sin x \csc x\]

OpenStudy (anonymous):

\[\frac{ \cot ^{2} x}{ 1+\csc x }+\sin x \csc x\]

OpenStudy (tkhunny):

\(\sin(x) = \dfrac{1}{\csc(x)}\)

OpenStudy (anonymous):

how would you combine \[\frac{ 1 }{ \csc } \frac{ \csc }{ 1}\]

OpenStudy (tkhunny):

1) "csc" doesn't mean anything. 2) How would you combine \(\dfrac{csc(x)}{csc(x)}\)

OpenStudy (anonymous):

\[\frac{ \csc x }{ \csc x } = 1 Right?\]

OpenStudy (tkhunny):

Very good.

OpenStudy (anonymous):

\[\frac{ \cot ^{2}x}{ 1+\csc x }+1\]

OpenStudy (anonymous):

\[\frac{ \csc ^{2}x-1 }{ 1+\csc x }\]

OpenStudy (tkhunny):

Multiply by \(\dfrac{1 - \csc(x)}{1 - \csc(x)}\). See if anything magical happens.